OpenStudy (anonymous):

a 20 foot piece of string is cut into two pieces so that the longer piece is 5 feet longer than twice the shorter piece. if the shorter piece is x feet long, what are the lengths of both pieces?

OpenStudy (campbell_st):

form an equation that can be solved. the shorter string is x the longer string is 2x+5 total length 20 then short + longer = 20 x + 2x + 5 = 20 3x + 5 =20 I'll let you solve the equation

OpenStudy (anonymous):

for what one?

OpenStudy (campbell_st):

you should get 5 feet for the shorter piece

OpenStudy (anonymous):

what about the longer one?

OpenStudy (campbell_st):

well its 2x + 5 long.... when x = 5